OpenStudy (anonymous):

what is 4^-2=1 over 16 in logarithmic form?

hartnn (hartnn):

\(\large 4^{-2} = \dfrac{1}{16} \) \(\large a^x =b \implies \log_a b=x\)

OpenStudy (anonymous):

can you explain how i would work this out

hartnn (hartnn):

sure, take log on both sides of the equation, what do u get ?

OpenStudy (anonymous):

so take the log of both sides

hartnn (hartnn):

yup

OpenStudy (anonymous):

well the log of 4^-2 is -1.204

hartnn (hartnn):

we need not evaluate it. we need the log form, so keep it as log 4^-2 now use the fact that log a^b = b log a then log 4^-2 = ... ?

OpenStudy (anonymous):

and the log of 1 over 16 is the same thing

hartnn (hartnn):

ofcourse, 4^-2 = 1/4^2 = 1/16 so their log will be same

OpenStudy (anonymous):

so my answer is the log of 4^-2=1/16 right

hartnn (hartnn):

the answer should be in log form let me write out the steps

hartnn (hartnn):